OpenStudy (anonymous):

2 -- % X70 = _____ 4 Select one of the options below as your answer:A. 168 B. 175 C. 28 D. 35

OpenStudy (freckles):

\[\%=\frac{1}{100}\] So we have \[\frac{2}{4} \cdot \frac{1}{100} \times 70 \]

OpenStudy (mertsj):

\[\frac{2}{4}percent \times70=.005\times70=.35\]

OpenStudy (mertsj):

So I'm thinking something is not posted correctly...either the problem or the answers